OpenStudy (secret-ninja):

Given the equation y - 4 = 3 over 4 (x + 8) in point slope form, identify the equation of the same line in standard form. -3 over 4x + y = 10 3x - 4y = -40 y = 3 over 4x + 12 y = 3 over 4x + 10

OpenStudy (imstuck):

The original equation is given as this:\[y-4=\frac{ 3 }{ 4 }(x+8)\]In standard form, the x and the y are together on one side of the equals sign with the everything else on the other side. So the first thing to do is to distribute the 3/4 into the parenthesis to get rid of them. That will give you this:

OpenStudy (imstuck):

\[y-4=\frac{ 3 }{4 }x+6\]

OpenStudy (imstuck):

Now get rid of the denominator in the 3/4 by multiplying everything by 4, like this:

OpenStudy (imstuck):

\[4y-16=3x+24\]

OpenStudy (imstuck):

Now move the 3x over to the side with the y, and move the -16 over to the side with the 24. Like this:\[4y-3x=40\]and to have it in Ax + By = C, change around the x and y to look like this:\[-3x+4y=40\]

OpenStudy (imstuck):

OR you could change the signs on everything to reflect what they say your answer should be, which is 3x-4y=-40

OpenStudy (secret-ninja):

So it would be B?

OpenStudy (imstuck):

Yep, it would be B.

OpenStudy (secret-ninja):

Thank you! I really think I over thought it lol It was simpler than I thought! Thanks!

OpenStudy (imstuck):

It usually is simpler than we think when we learn the simple way to do it, right?! : )

OpenStudy (imstuck):

TY for the medal!